# QUIDS: Query Intent Generation via Dual Space Modeling

### v2

@@ 3.3.1. Representation space modeling
(3) $\mathcal{L}_{encoder}=\mathcal{L}_{rel}+\mathcal{L}_{irrel}$
3.3.2. Disentangling space modeling
In the previous encoding phase, we obtain contextual feature representations for relevant and irrelevant documents given a query. In the decoder part, we want to make sure the decoder generates intent descriptions based on these information sources. Hence, we design a contrastive decoder by adding an additional cross attention so that it can attend to both relevant and irrelevant documents. To effectively disentangle the contrastive information learned by the model during the decoding process, we apply additional contrastive learning in disentangling space. As shown in Figure 2 (d), the model is designed to capture the contrast between relevant and irrelevant information spaces, allowing it to model both relevant and irrelevant intent spaces. This enables the generation of more precise and nuanced intent descriptions by focusing on the relevant intent space while minimizing attention on the irrelevant one.
@@ 3.3.2. Disentangling space modeling
(5) $h_{irrel,z}^{l}=MHAtt(h_{self,z}^{l},h_{i^{\prime}})$
@@
(6) $h_{combine,z}^{l}=h_{self,z}^{l}+h_{rel,z}^{l}-h_{irrel,z}^{l}$
(7) $P^{vocab}_{z}=Softmax(W(h_{combine,z}^{N}))$
where $W$ indicates a linear transformation. We optimize the model with the negative log likelihood (NLL) objective function to predict the target words:
(8) $\mathcal{L}_{NLL}=-\sum^{|y|}_{z=1}logP^{vocab}_{z}(y_{z})$
@@
(9) $\mathcal{L}_{decoder}=-log\frac{exp(cos(f_{c},f_{r})/\tau)}{\sum_{i^{\prime}\in\mathcal{B}}exp(cos(f_{c},f_{i^{\prime}})/\tau)}$
where $\tau$ is the temperature and $cos(\textperiodcentered,\textperiodcentered)$ defines cosine similarity.
@@
(10) $\mathcal{L}_{NLL}=\lambda_{0}\mathcal{L}_{NLL}+\lambda_{1}\mathcal{L}_{Encoder}+\lambda_{2}\mathcal{L}_{Decoder}$
@@
where $\lambda$ parameters control the balance between the three losses, with their total sum equal to 1.
4. Experimental Settings
4.1. Dataset
@@
Q2ID (Zhang et al., 2020) is a benchmark dataset for query-to-intent description, derived from existing TREC and SemEval collections: TREC is an ongoing series of benchmark activities in IR. Q2ID comprises the TREC Dynamic Domain track 2015, 2016, 2017 on dynamic, exploratory search and the TREC 2004 Robust Track on consistency of retrieval technology. These tracks together contribute to 510 entries in the Q2ID benchmark dataset. SemEval is a series of workshops on Semantic Evaluation. Q2ID comprises the English SemEval-2015 Task3 and SemEval-2016 Task3 on Community Question Answering, contributing 4878 entries to the Q2ID dataset.
For each query in the collections, documents with multi-graded relevance labels were converted into binary labels, indicating whether they are relevant or not to the query. In total, the dataset comprises 5,358 entries, each formatted as a quadruple: ¡query, relevant documents, irrelevant documents, intent description¿, with intent descriptions being human-written narratives that provide detailed descriptions of the search intent behind the queries. The length statistics are an average of 7.2 tokens per query and 45.5 tokens per intent description. We split the 5338 queries into training (5000), validation (100), and test (258) query sets according to the original split setting of the Q2ID benchmark dataset.
4.2. Baselines
Since query-focused summarization task and query intent generation task are closely aligned in their focus on user queries and the extraction of relevant information, we evaluate our model against state-of-the-art models from the Q2ID benchmark, QFS task, and additionally include LLM zero-shot and two-shot baselines.
  • Large Language Model (LLM) baseline: LLama3.1 (AI, 2024): We use the Llama3.1-8B instruction-tuned text-only model in both zero-shot and two-shot settings and conduct five experimental runs for each setting. For two-shot setting, we randomly using two different examples per run – one sourced from TREC and the other from SemEval.
  • Query Focused Summarization (QFS) baselines: RelReg (Vig et al., 2022) and RelRegTT (Vig et al., 2022): two-step approaches for QFS consisting of an score-and-rank extractor and an abstractor. The extractor is trained to predict ROUGE relevance scores and then the ranked results based on ROUGE are passed to the abstractor. SegEnc (Vig et al., 2022): an end-to-end approach tailored for handling longer input texts. SegEnc splits a long input into fixed-length overlapping segments and encodes them separately. The encoding sequences are concatenated so that the decoder can attend to all encoded segments jointly. Socratic (Pagnoni et al., 2023): an unsupervised, question-driven pretraining approach designed to tailor generic language models for controllable summarization tasks. Qontsum (Sotudeh and Goharian, 2023): an abstractive summarizer that applied Generative Information Retrieval (GIR) techniques. It builds on SegEnc by adding a segment scorer and contrastive learning modules.
We train the QFS baselines on the Q2ID dataset using the original code provided by the authors, except for Qontsum, which we independently reproduced.
4.3. Implementation Details
We build our method based on the BART-large (Lewis et al., 2020) model using Huggingface Transformers (Wolf et al., 2019). CrossRel attention and CrossIrrel attention layers in the decoder blocks share the same initial weights from the pre-trained BART-large model. For the IDNA step, we set the expected number of irrelevant documents per query to three in the training set. To ensure this, we augment 1,984 queries so that each query includes at least three irrelevant documents. We train the model for 10 epochs using the Adam optimizer with a learning rate of 0.0001. The final checkpoints are chosen based on the average ROUGE-{1, 2, L} scores against the ground truth query intent obtained from the validation set. During decoding, we set the maximum length to 256 tokens and apply beam search with a beam size of 4, using a no-repeat n-gram size of 3. We optimize $\lambda$ during training and fix on $(\lambda_{0}=0.2,\lambda_{1}=0.2,\lambda_{2}=0.6)$ to balance the three losses. All experiments are conducted five times, and we report the average results across these runs. The implementation details for the other baseline models can be found in Appendix A.1.
4.4. Evaluation Metrics
@@
We conduct three kinds of evaluations using different evaluator resources: automatic evaluation, LLM-based evaluation and human evaluation. For automatic evaluation, we report recall scores on ROUGE-{1, 2, L} following Zhang et al. (2020), along with BERTScore (Zhang et al., 2019), which assesses semantic and syntactic similarity beyond exact word matches. Additionally, we conduct a human evaluation study using 50 randomly selected test samples. Three PhD students in Computer Science scored intent descriptions from our model and the best baseline, without knowing which model produced them. They rated both models on four customized qualitative criteria, with scores ranging from 1 (worst) to 5 (best), and the final scores were averaged across the three annotators.
  • Fluency: to what extent the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions.
  • Factual Alignment: to what extent the generated query intent description is factually aligned with the ground truth intent.
  • Inclusion score: how well the generated query intent includes important details from the query and relevant documents.
  • Exclusion score: how well the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents.
However, conventional automatic evaluation metrics tend to show weak correlation with human judgments. Recent study (Fu et al., 2024; Liu et al., 2023b) propose using LLMs as evaluators for natural language generation (NLG) systems, scoring generated outputs based on generation probability rather than reference targets. Following the method of (Liu et al., 2023b), we use LLaMa3.1-8B [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ as instruction-tuned evaluators to assess the generated intent across four qualitative metrics. Specifically, we define the evaluation task and criteria, prompting the LLM to generate chain-of-thoughts (CoT) for the ‘Evaluation Steps’. For LLaMa3.1-8B, we use the output token probabilities from the LLMs to normalize the scores and take their weighted summation as the final results: $score=\sum_{i=1}^{n}p(s_{i})\times s_{i}$ where $S=\{s_{1},s_{2},...,s_{n}\}$ represents the predefined score set from the prompt, with a maximum value of 5 in our case. For the close-sourced GPT-4o, we sample 20 times to estimate the token probabilities. Example prompts are in Appendix B.
5. Experimental Results
Table 1. Performance between our model and baselines in terms of automatic evaluation (%). $\dagger$ indicates reported performance from previous work. ‘-’ means the result is inaccessible. * indicates the model outperforms the best baseline significantly with paired t-test at $p$-value $<$ 0.05 level. The best results are highlighted in bold, while the best baseline results are underlined.
@@ 5. Experimental Results
LLaMa3.1 zero-shot 29.28 7.42 20.90 57.26
LLaMa3.1 2-shot 32.75 9.54 24.34 57.89
@@
QUIDS_BART 34.47* 14.86* 26.77* 63.55*
5.1. Overall Results
We compare model performance between QUIDS and baselines in Table  1. The results show that: (1) Our model QUIDS outperforms all other baselines; (2) Our approach is compatible with both T5 and BART architectures, which both have a transformer-based encoder-decoder structure. This demonstrates the effectiveness of our design in the query intent generation task. Notably, BART-large outperforms T5-large despite having nearly half the model size, highlighting its efficiency in this task. (3) We implemented QFS models for the query intent generation task, which already outperform the baseline model, CtrsGen. However, our design further improves ROUGE-1 by 2.64 and ROUGE-L by 1.59 compared to the best QFS model, SegEnc; (4) In our experiments, the two-shot setting with the LLaMa3.1 8B model significantly outperforms the zero-shot setting in ROUGE scores, while showing only minor improvements in BERTScore. This suggests that without fine-tuning, intent generation may be lexically similar but semantically misaligned.
5.2. Ablation Study
Table 2. Ablation study of our QUIDS model with its variants under automatic evaluation (%).
Models RG-1 RG-2 RG-L BS
@@ 5.2. Ablation Study
We conduct an ablation study on the test set to assess the impact of various modules in our proposed approach under three conditions: (1) We remove the IDNA module to examine the effect of data preprocessing (w/o IDNA); (2) We remove contrastive learning from the encoder and (3) decoder separately to assess the importance of representation space modeling (w/o RSM) and disentangling space modeling (w/o DSM), respectively. Table 2 shows the results. Removing IDNA leads to a 1.16-point drop in BERTScore, suggesting that incorporating harder irrelevant documents helps the model capture relevant intent information more effectively. Removing contrastive learning from the decoder leads to a significant drop in all metrics, highlighting its importance in modeling a distinguishable intent space and generating accurate intent descriptions. Finally, removing contrastive learning from the encoder results in a slight decrease, indicating that representation space modeling aids the model in distinguishing relevant from irrelevant information. Due to the similar architecture and experimental performance, we omit the T5 model results.
5.3. Human and LLM Evaluation
We analyze the quality of the generated intents by our model QUIDS and the best baseline model SegEnc with human evaluation and LLM-based approaches. Human inter-annotator agreement is assessed with weighted Cohen’s kappa, revealing fair to moderate consistency across the four metrics. As shown in table 3, QUIDS consistently receives higher scores from all three evaluators, except for Factual Alignment, where human evaluators prefer SegEnc’s outputs.
Table 3. Comparison of human evaluation and LLM evalaution in terms of Fluency (Fluen.), Factual Alignment (Align.), Inclusion score (Inclu.) and Exclusion score (Exclus.).
Method Model Fluen. Align. Inclu. Exluc.
Human SegEnc 4.71 3.70 3.88 4.71
QUIDS 4.73 3.61 3.97 4.73
@@ 5.3. Human and LLM Evaluation
Figure 3. Distribution of human and LLM evaluation sores on four qualitative metrics.
[…]
Figure 4. (a) Visualizes the token-level decoder cross-attention weights for a relevant document snippet when generating the token ‘impacted’ in Intent 1, without an irrelevant document provided. (b) for relevant document and irrelevant document when generating the token ‘effects’ in Intent 2, with an irrelevant document provided. Deeper color indicates a higher value.
[…]
To examine the performance differences between humans and LLMs in this task, we present boxplots in Figure 3 that display the score distributions for each metric. Additionally, we calculate the Spearman and Kendall-Tau correlations between the evaluations from LLMs and human judgments, as shown in Table 5 in Appendix A.2. From this analysis, we derive the following insights:
  • Humans tend to assign higher scores than GPT-4o and LLaMa 3.1 across all metrics, particularly for Fluency and Exclusion, where scores nearly reach 5. The variability in fluency scores suggests humans are more tolerant of less fluent text.
  • Human evaluations exhibit the narrowest range, while GPT-4o shows a similar range but with more critical scores. As supported by the correlation table in Appendix A.2, GPT-4o aligns more closely with human evaluations, particularly in Factual Alignment. However, a difference in model preferences for this metric prompts further analysis of sub-dataset differences in Section 5.4.
  • For the metrics Factual Alignment and Inclusion, all methods tend to exhibit a wider distribution with generally lower scores. This suggests variability in the assessments of these two metrics, potentially indicating inconsistency in how both human and LLM evaluators interpret the scoring criteria.
5.4. Analysis of Intent Types
For analysis of the effect of intent type, we classify the queries according to their underlying search intent into two categories: (1) Informational Intent: Natural language questions seeking detailed information or solutions, typically longer and contextual. Queries from the SemEval dataset fall under this category. (2) Exploratory Intent: Term-based queries aimed at broad exploration with minimal context or structure. Queries from the TREC datasets are categorized here. In the 258 test samples, there are 20 queries with exploratory intents and 238 with informational intents. As shown in Table 4, queries with exploratory intents substantially outperform those with informational intents, achieving 60% higher ROUGE-2 scores and 20% higher BERTScores. This indicates that our model is better suited for exploratory queries. This finding contrasts with the results of (Zhang et al., 2020), where the CtrsGen model performed slightly better on the informational SemEval queries than on the exploratory TREC queries. A potential explanation is that the backbone language model used in our approach more freely generates text than the GRU model used in (Zhang et al., 2020), particularly when reconstructing complex scenarios for informational intents. This is an aspect that makes our approach more suitable for exploratory search rather than informative search. We additionally conduct a breakdown analysis by intent types for the human and LLM results in Appendix A.3.
Table 4. Comparison of automatic evaluation on our model QUIDS for different intent types.
@@ 5.4. Analysis of Intent Types
5.5. Analysis of Document Length
In real-world scenarios, relevant documents associated with a query can vary in length. To assess the impact of document length on the generated output, we categorize the documents into three groups based on the BART-large input limit (1024 tokens): documents shorter than 512 tokens are classified as short documents; those ranging from 512 to 1024 tokens are termed medium-sized documents; and documents longer than 1024 tokens, which the model will truncate, are labeled as long documents. In our test dataset, there are 204 short, 25 medium, and 29 long documents. The overall differences in performance among all the input length categories are minimal across all metrics (also see Figure 6 in Appendix A.4). This indicates that 1024 tokens provide sufficient information for our model to effectively capture the underlying query intent. This may explain why QFS approaches, particularly those designed to handle lengthy documents, do not demonstrate significant advantages in the query intent generation task.
5.6. Case Study
To intuitively understand how irrelevant documents influence our model, we compare the generated intents for a query ‘Freon-12’ from test set with and without providing an irrelevant document. Figure 4 illustrates the token-level decoder cross-attention weights for both a relevant document snippet, and an irrelevant document snippet when provided. As shown in Figure 4 (a), without an irrelevant document provided, when generating the word ‘impacted’, the model’s focus is centered on the economic effects on cars, as indicated by the mention of ‘price’ in Intent 1. With an irrelevant document provided in Figure 4 (b), when generating the word ‘effects’ in Intent 2, similar attentions can be detected on ‘prices’, ‘\$’, ‘costs’, ‘auto’ for both relevant and irrelevant documents. However, as shown in Generated Intent 2, our model now treats tokens related to prices and cars in relevant documents as irrelevant. This demonstrates that out model effectively generates intent descriptions that excludes this unrelated information.
6. Conclusion

### v3

@@ 3.3.1 Representation Space Modeling (RSM)
$\mathcal{L}_{encoder}=\mathcal{L}_{rel}+\mathcal{L}_{irrel}$ (3)
3.3.2 Disentangling Space Modeling (DSM)
In decoding, we aim to generate intent descriptions based on the encoded relevant and irrelevant document features. To achieve this, we design a contrastive decoder with an added cross-attention layer that attends to both sources. To further disentangle relevant from irrelevant information, we apply contrastive learning in a separate disentangling space. As shown in Figure 2(d), this helps the model focus on relevant intent while minimizing influence from irrelevant content, enabling more precise and nuanced intent generation.
@@ 3.3.2 Disentangling Space Modeling (DSM)
$h_{irrel,z}^{l}=MHAtt(h_{self,z}^{l},h_{i^{\prime}})$ (5)
@@
$h_{combine,z}^{l}=h_{self,z}^{l}+h_{rel,z}^{l}-h_{irrel,z}^{l}$ (6)
$P^{vocab}_{z}=Softmax(W(h_{combine,z}^{N}))$ (7)
where $W$ indicates a linear transformation. We optimize the model with the negative log likelihood (NLL) objective to predict the target words:
$\mathcal{L}_{NLL}=-\textstyle\sum^{|y|}_{z=1}logP^{vocab}_{z}(y_{z})$ (8)
@@
$\mathcal{L}_{decoder}=-log\frac{exp(cos(f_{c},f_{r})/\tau)}{\sum_{i^{\prime}\in\mathcal{B}}exp(cos(f_{c},f_{i^{\prime}})/\tau)}$ (9)
where $\tau$ is the temperature and $cos(\cdot,\cdot)$ defines cosine similarity.
@@
$\mathcal{L}_{NLL}=\lambda_{0}\mathcal{L}_{NLL}+\lambda_{1}\mathcal{L}_{Encoder}+\lambda_{2}\mathcal{L}_{Decoder}$ (10)
@@
where the $\lambda$ parameters control the balance between the three losses, with their total sum equal to 1.
4 Experimental Settings
4.1 Data
@@
We conduct experiments on the Q2ID dataset (Zhang et al., 2020), a benchmark for query-to-intent description derived from existing TREC and SemEval collections. Specifically, it comprises: TREC: Including the Dynamic Domain tracks (2015–2017) and the 2004 Robust Track, which focus on dynamic, exploratory search and consistency of retrieval technology. SemEval: Including the English SemEval-2015 and SemEval-2016 Task 3 tracks on Community Question Answering. Q2ID contains a total of 5,358 entries. Each entry is structured as a quadruple: <query, relevant documents, irrelevant documents, intent description>, where the intent descriptions are human-written narratives. The statistics and more details are provided in Appendix B.1.
4.2 Baselines
To reflect the shared focus on user queries and the extraction of relevant content, we compare our model with baselines from four categories: (i) Pretrained Seq2Seq Models: We fine-tune T5-large (Raffel et al., 2020b) and BART-large (Lewis et al., 2020) on the Q2ID dataset. BART also serves as the backbone of our QUIDS model. (ii) Q2ID Baseline: CtrsGen (Zhang et al., 2020) leverages contrastive generation using a bi-GRU encoder and contrast-weighted attention mechanism. (iii) LLM Baseline: We evaluate a instruction-tuned model LLaMA3.1-8B-Instruct (AI, 2024) and a reasoning model OpenAI o3 (OpenAI, 2025) in zero-shot and two-shot settings. For the two-shot settings, examples are drawn from TREC and SemEval. (iv) QFS Baselines: We include extractive-abstractive models RelReg, RelRegTT (Vig et al., 2022), the segment-based model SegEnc (Vig et al., 2022), the question-driven Socratic (Pagnoni et al., 2023), and the contrast-enhanced Qontsum (Sotudeh and Goharian, 2023). Detailed descriptions, training configurations, and reproduction settings for baselines are in Appendix B.4 and B.5.
4.3 Implementation Details
Our method is implemented based on the BART-large model (Lewis et al., 2020) using the Huggingface Transformers library (Wolf et al., 2019). CrossRel and CrossIrrel attention layers in the decoder are initialized with pre-trained BART weights. We optimize the weighted training loss using coefficients $(\lambda_{0}=0.2,\lambda_{1}=0.2,\lambda_{2}=0.6)$ to balance multiple objectives (see Appendix B.2). The model is trained with the Adam optimizer, and the final checkpoint is selected based on average ROUGE-{1, 2, L} scores on the validation set. We provide additional training details in Appendix B.3.
4.4 Evaluation Metrics
@@
We conduct three types of evaluations using different evaluator resources: automatic evaluation, LLM-based evaluation and human evaluation. For automatic evaluation, we report recall scores on ROUGE-{1, 2, L} following Zhang et al. (2020), along with BERTScore (Zhang et al., 2019), which assesses semantic and syntactic similarity beyond exact word matches. We also conduct a human evaluation study using 50 (Sotudeh and Goharian, 2023) randomly selected test samples. Five PhD students in Computer Science scored intent descriptions from our model and the best baseline, without knowing which model produced them. They rated both models on four customized qualitative criteria with scores ranging from 1 (worst) to 5 (best). Four criteria are: (1) Fluency: to what extent the generated query intent description reads naturally, understandably, and without noticeable errors or disruptions. (2) Factual Alignment: to what extent the generated query intent description is factually aligned with the ground truth intent. (3) Inclusion score: how well the generated query intent includes important details from the query and relevant documents. (4) Exclusion score: how well the generated query intent description excludes information present in the irrelevant documents that is not relevant to the query and relevant documents. Inspired by Liu et al. (2023b), we also adopt LLM-based evaluation (LLaMa3.1-8B-Instruct and GPT-4o) by prompting instruction-tuned models to assess generations across the same four qualitative metrics. Details on the prompt formats, and scoring computation are provided in Appendix F.
@@
LLaMa3.1 (0) 29.28 7.42 20.90 57.26
LLaMa3.1 (2) 32.75 9.54 24.34 57.89
o3 (0) 34.91 6.84 23.93 59.19
o3 (2) 33.15 6.28 22.81 59.88
@@
QUIDS_BART 34.47 14.86* 26.77* 63.55*
Table 1: Performance between our model and baselines in terms of automatic evaluation (%). $\dagger$ indicates reported performance from previous work. ‘-’ means the result is inaccessible. * indicates the model outperforms the best baseline significantly with paired t-test at $p$-value << 0.05 level. Results are averaged over 5 random seeds. The best results are highlighted in bold, while the best baseline results are underlined.
Model RG-1 RG-2 RG-L BS
@@
Table 2: Ablation study of our QUIDS model with its variants under automatic evaluation (%).
Figure 3: Distribution of human and LLM evaluation sores on four qualitative metrics.
5 Experimental Results
5.1 Overall Results
We compare model performance between QUIDS and baselines in Table 1. The results show that: (1) QUIDS outperforms all baselines except o3 on RG-1, including those with larger model sizes, indicating that the gains stem primarily from our dual-space modeling design rather than model size or pretraining knowledge; (2) Our approach is compatible with both T5 and BART architectures. Notably, BART-large outperforms T5-large despite having nearly half the model size; (3) The QFS models that we implemented for the query intent generation task outperform the Q2ID-specific baseline, CtrsGen. QUIDS further significantly outperforms the best QFS model, SegEnc; (4) The two-shot setting with the LLaMa3.1-8B-Instruct model significantly outperforms the zero-shot setting in ROUGE scores, while showing only minor improvements in BERTScore. This suggests that without fine-tuning, generated intents may be lexically similar but semantically misaligned; (5) While o3 shows strong reasoning and generates richer descriptions, it does not outperform our model on ROUGE and BERTScore. We observed that o3 often paraphrases intent with different wording, leading to lower ROUGE scores. Its longer outputs and inclusion of detailed justifications or summaries may dilute the concise intent signal. This suggests that strong reasoning alone may not align well with the goal of generating concise, system-inferred intent.
5.2 Ablation Study
We perform an ablation study based on the BART-Large model to evaluate the contribution of key components in our approach under three settings: (1) without the IDNA module (w/o IDNA), (2) without contrastive learning in the encoder (w/o RSM), and (3) without contrastive learning in the decoder (w/o DSM). Results are shown in Table 2. Excluding contrastive learning from the decoder leads to the largest performance drop, underscoring its role in modeling a discriminative intent space. Removing it from the encoder results in a smaller decline, suggesting that representation space modeling still contributes to relevance awareness.
Figure 4: Boxplot of human and LLM evaluation scores on 4 metrics of our model on different intent types.
5.3 Human and LLM Evaluation
We assess the quality of generated intents from QUIDS and the best baseline SegEnc using both human and LLM-based evaluations. Inter-annotator agreement, measured by weighted Cohen’s $\kappa$, and LLM-human correlations, measured by Spearman and Kendall $\tau$, indicate fair to moderate consistency across metrics (Table 7, Appendix C). As shown in Table 3, QUIDS outperforms SegEnc on all metrics except Factual Alignment, where humans prefer SegEnc. Further analysis (subsection 5.4.1) shows this stems from the dominance of informational queries, on which SegEnc performs better. QUIDS, by contrast, performs better on exploratory queries. Figure 3 further illustrates score distributions, revealing three key insights:
Method Model Fluen. Align. Inclu. Exclu.
Human SegEnc 4.75 3.90 3.94 4.77
QUIDS 4.80 3.80 4.06 4.80
@@ 5.3 Human and LLM Evaluation
Table 3: Comparison of human evaluation and LLM evalaution in terms of Fluency, Factual Alignment, Inclusion score and Exclusion score.
(1) Human scores are generally higher than LLM scores, especially for Fluency and Exclusion. The larger variability for fluency scores suggests humans may tolerate minor fluency issues. (2) Human evaluations show broader and lower score distributions for Factual Alignment and Inclusion, aligning more with LLaMa3.1 (Table 7). In contrast, they mirror GPT-4o’s narrower distribution for Fluency and Exclusion, where correlation is higher. This suggests that evaluators differ in how they assess each metric. (3) Fluency and Factual Alignment show stronger alignment between LLM and human evaluations, likely due to being less context-dependent. In contrast, Inclusion and Exclusion scores exhibit weaker correlations, indicating inconsistencies in evaluating context-sensitive criteria.
@@
Table 4: Comparison of automatic evaluation on our model for different intent types.
Figure 5: Case study indicating the role of contrastive examples in the decoder stage. Token-level decoder cross-attention weights are shown for a generated intent token (red) are shown with (a) and without (b) an irrelevant document in the model input. Deeper color indicates a higher value.
5.4 In-depth Analysis
5.4.1 Analysis of Intent Types
We classify the queries according to their underlying search intent into two categories: (1) Informational Intent: Natural language questions seeking detailed information or solutions, typically longer and contextual. Queries from the SemEval dataset fall under this category. (2) Exploratory Intent: Term-based queries aimed at broad exploration with minimal context or structure. Queries from the TREC datasets are categorized here. Our automatic and human evaluation by intent type indicates that QUIDS is indeed more successful for exploratory tasks than for informational intent tasks. Results are shown in Table 4 and Figure 4. See Appendix D for detailed analysis.
Figure 6: Robustness analysis on (a) input length and (b) IDNA threshold.
5.4.2 Analysis of Document Length
In real-world scenarios, relevant documents associated with a query can vary in length. Documents over 1024 tokens are truncated in our architecture. To assess robustness to input length, we group test documents into short (<512 tokens), medium (512–1024), and long (>1024). In the test dataset, 204 are short, 25 medium, and 29 long. Figure 6(a) shows that the overall differences in performance among all the input length categories are minimal across all metrics, suggesting that 1024-token inputs are sufficient for inferring the underlying query intent. This may explain why long-document QFS models offer limited gains in this task.
5.4.3 Sensitivity Analysis of IDNA threshold
Intent-driven negative augmentation (IDNA) selects irrelevant documents with high similarity to relevant ones as hard negatives for contrastive learning. We noticed that most document similarity scores fall between 0.6 and 0.9, making the method robust even with lower thresholds (e.g., <0.6). Performance improves as the threshold increases within this range, with 0.8 yielding the best results, which we use in our experiments.
5.5 Case Study
Figure 5 illustrates how the model uses cross-attention in the decoder stage to identify irrelevant semantics from a low-ranked document. When generating ‘impacted’ without an irrelevant document (Figure 5(a)), the model focuses on economic effects on cars, indicated by ‘price’ in Intent 1. With an irrelevant document in Figure 5(b), while similar economic attentions are observed across both relevant and irrelevant documents when generating the word ‘effects’ in Intent 2, the model successfully identifies tokens related to prices and cars in relevant documents as irrelevant. This demonstrates the model’s ability to filter out irrelevant content using contrastive learning in the decoder. We include another failure case study in Appendix E.
6 Conclusions